Grade 11Physical Chemistry

If isotopic distribution of c-12 & c-14 is 98% &2% respectively calculate 5he no. Of c-14 atoms in 12 gm of carbon

Kanak Pitalia
9 Years agoGrade 11

2 Answers

Vikas TU
9 Years ago
12 gms means 1 mole of carbon atoms in total => Na.
Let x be the C-12 no. of atoms then C-14 would have (Na – x)
Therfore,
((98*x + (Na – x)*2)/100) = Na
98x + 2Na – 2x = 100Na
96x = 98Na
x = 98Na/96
and for C-14 it would be =>  98Na/96 – Na=> 2Na/96
Sunny Kumar
8 Years ago
This ans is wrong
According to this 
C¹⁴ molecules=Na-x=Na-98Na/96=Negative
So correct ans is 1.032*10²²